# From 2D Images to 3D Model: Weakly Supervised Multi-View Face Reconstruction with Deep Fusion

### 04 Jan 2024 v3

@@ 3.4 Parametric Regression
$\begin{array}[]{l}\mathbf{S}=\mathbf{S}(\boldsymbol{\alpha},\boldsymbol{\beta})=\overline{\mathbf{S}}+\mathbf{B}_{id}\boldsymbol{\alpha}+\mathbf{B}_{exp}\boldsymbol{\beta},\\ \mathbf{T}=\mathbf{T}(\boldsymbol{\gamma})=\overline{\mathbf{T}}+\mathbf{B}_{t}\boldsymbol{\gamma},\end{array}$ (1)
where $\overline{\mathbf{S}}$ and $\overline{\mathbf{T}}$ are the average face shape and texture. $\mathbf{B}_{id}$, $\mathbf{B}_{exp}$, $\mathbf{B}_{t}$ are the PCA bases of identity, expression, and texture respectively. $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, and $\boldsymbol{\gamma}$ are the parameter vectors that the network needs to regress ($\alpha,\beta\in R^{80}$ and $\gamma\in R^{64}$). By adjusting these three vectors, the shape, expression and texture of the 3D face can be changed. In order to compare with MGCNet Shang et al. (2020) and Deep3DFace Deng et al. (2019) fairly, we use the same face model. BFM Paysan et al. (2009) is adopted for $\overline{\mathbf{S}},\mathbf{B}_{id},\overline{\mathbf{T}}$, and $\mathbf{B}_{t}$. $\mathbf{B}_{exp}$ Guo et al. (2018) is built based on Facewarehouse Cao et al. (2013).
Pose Parameters are used to adjust the angle and position of the 3D face in the camera coordinate system. We exploit the differentiable perspective rendering Ravi et al. (2020) to render the 3D face back to 2D. When the camera coordinates are fixed, we could change the size and angle of the rendered 2D face by adjusting the position of the 3D face in the camera coordinate system. Meanwhile, the position of the 3D face in the camera coordinate system can be determined by predicting the rotation angle and translation in each axis. In order to enhance the geometric constraints of the multi-view reconstruction, we respectively predict the pose of the 3D faces in the multi-view Dai et al. (2023), instead of only predicting the pose of one perspective to render 2D images.
@@ 4.1 Photo Loss
Photo loss is often used in weakly supervised face reconstruction tasks Thies et al. (2016); Tewari et al. (2018); Deng et al. (2019); Shang et al. (2020). Distinct with the traditional methods, we impose a weight for each pixel according to the facial features. The weight map is learned by the face mask $\mathcal{M}$ of the original image $I$. In order to enhance the robustness of the weight map, we dilate $\mathcal{M}$ with 20 pixel as $\mathcal{M}_{d}$, as shown in Fig. 4. The dilated image is divided into three levels to be the weight map. In weight map, facial features are marked as 254, the rest of the facial area is marked as 128, and the background is marked as 32. The multi-view photo loss can be expressed as:
@@ 4.2 Mask Loss
Photo loss focuses on the pixel difference between two pictures. It is difficult to constrain the size of the facial feature area in the two pictures. For example, the nose color is very similar to that of the cheeks, thereby leading to difficulties for the photo loss to notice the boundary line between them. For this reason, we introduce mask loss to narrow the facial features of the input image and the rendered image. The division and labeling of the facial features are shown in Fig. 2. According to BiSeNet Yu et al. (2018) we adopt cross entropy to calculate the mask loss $L_{m}$:
@@ 4.3 Landmark Loss
We also adopt 2D landmarks and 3D landmarks for weakly supervised training. We use 3D face alignment method Bulat and Tzimiropoulos (2017) to generate 68 landmarks $\left\{l_{n}\right\}$ as the ground-truth. Then the corresponding points in the predicted 3D face point cloud are projected to 2D as predicted 2D landmarks $\left\{l_{n}^{\prime}\right\}$. Then the multi-view 2D landmark loss can be calculated:
@@
2D landmarks are still insufficient for the reconstruction of 3D face shapes. In order to obtain better reconstruction effect, we select 101 3D landmarks $\left\{q_{n}^{\prime}\right\}$ to impose a weak constraint on the shape of the 3D face. According to the 3DMM index, 101 predicted landmarks $\left\{q_{n}\right\}$ can be found. Then, we select 7 points $\left\{a_{n}^{\prime}\right\}$ and $\left\{a_{n}\right\}$ in $\left\{q_{n}^{\prime}\right\}$ and $\left\{q_{n}\right\}$ respectively as alignment points to calculate the alignment parameters of $\left\{q_{n}^{\prime}\right\}$ and $\left\{q_{n}\right\}$. The alignment parameters include: scale $\mathbf{s}$, rotation $\mathbf{R}$ and translation $\mathbf{t}$. These parameters can be obtained by the following optimization equation Sanyal et al. (2019):
$Optim(\mathbf{s},\mathbf{R},\mathbf{t})=\min_{\mathbf{s},\mathbf{R},\mathbf{t}}\sum_{i}\left\|\mathbf{a}_{i}^{\prime}-\mathbf{s}\left(\mathbf{R}\cdot\mathbf{a}_{i}+\mathbf{t}\right)\right\|_{2}.$ (5)
@@ 4.4 Regularization Loss
To suppress the generation of distorted faces, we add the regularization loss which is commonly-used in face reconstruction task Thies et al. (2016); Tewari et al. (2018); Deng et al. (2019); Shang et al. (2020):
@@
where $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, and $\boldsymbol{\gamma}$ are 3DMM parameter vectors that the network predicted. $\omega_{\alpha}$, $\omega_{\beta}$ and $\omega_{\gamma}$ are the weights for 3DMM parameter vectors. Following Deep3DFace Deng et al. (2019), we set them to $1$, $0.8$ and $0.017$ with fine tuning.
@@ 4.5 Overall Loss
Note that 3D landmarks is not a must for our method. For general weakly supervised methods, we do not incorporate $L_{l\_3d}$ for comparison Deng et al. (2019); Shang et al. (2020). Since semi-supervised methods HRN Lei et al. (2023) adopts the entire 3D scans to attain training label, we add $L_{l\_3d}$ for comparative analysis.
@@ 5.1 Implementation Details
Dataset. Following BiseNet Yu et al. (2018), we adopt CelebAMask-HQ Lee et al. (2020) to pretrain Face Parse Net. Furthermore, Pixel-Face Lyu et al. (2020) and Multi-PIE Gross et al. (2008) are introduced to provide 3D landmarks label and multi-view faces. Pixel-Face contains $855$ subjects ranging in age from $18$ to $80$ years old. Each subject has $7$ or $23$ samples of different expressions. Pixel-Face has 3D mesh file of each sample as ground-truth but not 3DMM parameters or angle of multi-view images. In the experiment, the training/test split was set to $0.8$. For Multi-PIE, we leverage the version provided by TP-GAN Huang et al. (2017), which will reduce a lot of data preprocessing work. This dataset contains about 140k multi-view images and does not provide 3D mesh and landmarks. We utilize the method of STNet Chen et al. (2016) to detect and align faces. Then each input image is cropped and resized to 224$\times$224.
Network. Our network structure is shown in Fig. 1 and described in the methodology section. Based on the pre-trained BiseNet Yu et al. (2018) with frozen weights, the face parse network is located in the beginning and end of the network. We implement DF-MVR by pytorch Paszke et al. (2019). It is trained on 2 NVIDIA RTX 3090 GPUs card with a batch size of 16 for 512 epochs. We adopt Adam as the gradient descent method. The initial learning rate is set to 0.0001 which decays with the cosine anneal schedule. The average inference time of DF-MVR is 244ms.
Evaluation Metric. Following the previous works, RMSE (mm) Wu et al. (2019); Deng et al. (2019); Shang et al. (2020) is used to compute point-to-plane L2 distance between predict 3D scans and ground-truth 3D scans. Concretely, the front face area is cropped for metrics calculation instead of using a complete BFM model Sanyal et al. (2019); Deng et al. (2019); Shang et al. (2020). Before calculating point-to-plane L2 distance, the predicted 3D scans need to be registered with ground-truth 3D scans. Also, we take the ICP registration method Li et al. (2017), the same as Deep3DFace Deng et al. (2019).
@@ 5.2 Comparison to SOTAs
We compare our method with the existing weakly supervised MVRs on both Pixel-Face and Bosphorus datasets. More specifically, only the four methods with codes can be found in the literature related to multi-view weakly supervised 3D face reconstruction. MGCNet Shang et al. (2020) takes multiple images for training, and then uses one single image for testing. We select the best results among the three images for display. Deep3DFace Deng et al. (2019) does not release their source codes of its scoring network. Therefore, we use their codes to train/test on Pixel-Face to select the best results among the three images. Moreover, MVFR Xiao et al. (2022) necessitates scale_RT data, rendering it currently effective only for the Facescape dataset Yang et al. (2020); Zhu et al. (2023). HRN Lei et al. (2023) does not release its training code, but it provides a pretrained model based on numerous data. To this end, we use its pretrained model and source code for multi-view face reconstruction. Given the scarcity of open sources for this task, we release our full codes (training & test) and pretrained model in GitHub [^4] https://github.com/weiguangzhao/DF_MVR.
@@
Deep3DFace Deng et al. (2019) Pixel-Face 1.6641
MGCNet Shang et al. (2020) Pixel-Face 1.8877
@@
${}^{\dagger}$ Lei et al. (2023) Pixel-Face 1.4061
@@
Deep3DFace Deng et al. (2019) Bosphorus 1.4777
MGCNet Shang et al. (2020) Bosphorus 1.4418
Ours Bosphorus 1.3994
Comparsion on Pixel-Face. The quantitative results of the comparison are shown in Table 1. Since MGCNet Shang et al. (2020) and Deep3DFace Deng et al. (2019) did not use 3D landmarks, to be fair, we also provide the results of our model without using 3D landmarks for comparison. As observed, Our model (without 3D landmarks) shows $5.2\%$ improvement compared to the existing methods. While HRN Lei et al. (2023) adopts the complete 3D ground-truth scan to attain deformation map and displacement map, we only take very light annotation, 101 3D landmarks to constraint the face reconstruction. ${}^{\dagger}$ shows very competitive results with fewer 3D annotation, compared to ${}^{\dagger}$.
@@
Shang et al. (2020) Pixel-Face 1.8877
Deng et al. (2019) Pixel-Face 1.6641
@@ 5.4 Ablation Study
In order to verify the effectiveness of Tri-Unet and the mask mechanism we designed, we perform more ablation experiments as shown in Table 3. The mean and standard deviation of RMSE are again used as the evaluation metric. First, from v1, v2, v6, it can be found that the multi-view feature fusion network we designed is superior to traditional CNN and Unet in this task. Then, the results of v2 and v6 hint that the multi-layer feature interaction in the feature extraction stage is better than the direct concatenation of features at the end. To be fair, we set the number of layers of RedNet and ResNet to 50. Through the RMSE of v3 and v6, it is clear that RedNet performs better than ResNet in this task. For v4, we not only remove the mask loss but also the face mask $\mathbf{I}_{A}$, $\mathbf{I}_{B}$ and $\mathbf{I}_{C}$, which are concatenated to the original image. By comparing v5 and v6, we can see that the face mask mechanism promotes the network to generate a higher-precision model. Moreover, we verify the face mask and mask loss separately. Mask loss has a much smaller effect on the accuracy of the model than face mask. While the qualitative result of mask loss is positive for the adjustment of the model during training, we retain this part for reference. Finally, we remove $L_{lan\_3d}$, which means that our model can be trained with only three multi-view images without any 3D label (as denoted as v5). The result also shows that our model is accurate and stable.
@@
v3 Tri-Unet ResNet Y Y 1.5206
v4 Tri-Unet RedNet N Y 1.4656
v5 Tri-Unet RedNet Y N 1.5770
v6 Tri-Unet RedNet Y Y 1.4040
@@ 5.5 Robustness Analysis
Shang et al. (2020) 1.4419 1.4412 1.4413 1.4418 1.4416
Deng et al. (2019) 1.5030 1.4975 1.4924 1.4777 1.4905
@@
We also validate the angle robustness of our model visually. These results are plotted in Fig. 8(a). Detailed error maps are introduced to display the error for each facial area directly. Take the second sample of Fig. 8(b) for a typical example where we fix $\mathbf{I}_{B}$ and $\mathbf{I}_{c}$ while changing the face angle of $\mathbf{I}_{A}$ from $10^{\circ}$ to $45^{\circ}$. Obviously, the error maps generated from these angles do not change much. Furthermore, we provide the error maps of Deep3DFace Deng et al. (2019) for comparison. At different angles, DF-MVR reconstructs the face with less error, especially in critical areas (e.g. eye, brow, nose and mouth).

### 22 Jan 2024 v4

@@ 3.4 Parametric Regression
$\begin{array}[]{l}\mathbf{S}=\mathbf{S}(\boldsymbol{\alpha},\boldsymbol{\beta}=\overline{\mathbf{S}}+\mathbf{B}_{id}\boldsymbol{\alpha}+\mathbf{B}_{exp}\boldsymbol{\beta},\\ \mathbf{T}=\mathbf{T}(\boldsymbol{\gamma}=\overline{\mathbf{T}}+\mathbf{B}_{t}\boldsymbol{\gamma},\end{array}$ (1)
where $\overline{\mathbf{S}}$ and $\overline{\mathbf{T}}$ are the average face shape and texture. $\mathbf{B}_{id}$, $\mathbf{B}_{exp}$, $\mathbf{B}_{t}$ are the PCA bases of identity, expression, and texture respectively. $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, and $\boldsymbol{\gamma}$ are the parameter vectors that the network needs to regress ($\alpha,\beta\in R^{80}$ and $\gamma\in R^{64}$). By adjusting these three vectors, the shape, expression and texture of the 3D face can be changed. In order to compare with MGCNet (Shang et al., 2020) and Deep3DFace (Deng et al., 2019) fairly, we use the same face model. BFM (Paysan et al., 2009) is adopted for $\overline{\mathbf{S}},\mathbf{B}_{id},\overline{\mathbf{T}}$, and $\mathbf{B}_{t}$. $\mathbf{B}_{exp}$ (Guo et al., 2018) is built based on Facewarehouse (Cao et al., 2013).
Pose Parameters are used to adjust the angle and position of the 3D face in the camera coordinate system. We exploit the differentiable perspective rendering (Ravi et al., 2020) to render the 3D face back to 2D. When the camera coordinates are fixed, we could change the size and angle of the rendered 2D face by adjusting the position of the 3D face in the camera coordinate system. Meanwhile, the position of the 3D face in the camera coordinate system can be determined by predicting the rotation angle and translation in each axis. In order to enhance the geometric constraints of the multi-view reconstruction, we respectively predict the pose of the 3D faces in the multi-view (Dai et al., 2023), instead of only predicting the pose of one perspective to render 2D images.
@@ 4.1 Photo Loss
Photo loss is often used in weakly supervised face reconstruction tasks (Thies et al., 2016; Tewari et al., 2018; Deng et al., 2019; Shang et al., 2020). Distinct with the traditional methods, we impose a weight for each pixel according to the facial features. The weight map is learned by the face mask $\mathcal{M}$ of the original image $I$. In order to enhance the robustness of the weight map, we dilate $\mathcal{M}$ with 20 pixel as $\mathcal{M}_{d}$, as shown in Fig. 4. The dilated image is divided into three levels to be the weight map. In weight map, facial features are marked as 254, the rest of the facial area is marked as 128, and the background is marked as 32. The multi-view photo loss can be expressed as:
@@ 4.2 Mask Loss
Photo loss focuses on the pixel difference between two pictures. It is difficult to constrain the size of the facial feature area in the two pictures. For example, the nose color is very similar to that of the cheeks, thereby leading to difficulties for the photo loss to notice the boundary line between them. For this reason, we introduce mask loss to narrow the facial features of the input image and the rendered image. The division and labeling of the facial features are shown in Fig. 3. According to BiSeNet (Yu et al., 2018) we adopt cross entropy to calculate the mask loss $L_{m}$:
@@ 4.3 Landmark Loss
We also adopt 2D landmarks and 3D landmarks for weakly supervised training. We use 3D face alignment method (Bulat and Tzimiropoulos, 2017) to generate 68 landmarks $\left\{l_{n}\right\}$ as the ground-truth. Then the corresponding points in the predicted 3D face point cloud are projected to 2D as predicted 2D landmarks $\left\{l_{n}^{\prime}\right\}$. Then the multi-view 2D landmark loss can be calculated:
@@
2D landmarks are still insufficient for the reconstruction of 3D face shapes. In order to obtain better reconstruction effect, we select 101 3D landmarks $\left\{q_{n}^{\prime}\right\}$ to impose a weak constraint on the shape of the 3D face. According to the 3DMM index, 101 predicted landmarks $\left\{q_{n}\right\}$ can be found. Then, we select 7 points $\left\{a_{n}^{\prime}\right\}$ and $\left\{a_{n}\right\}$ in $\left\{q_{n}^{\prime}\right\}$ and $\left\{q_{n}\right\}$ respectively as alignment points to calculate the alignment parameters of $\left\{q_{n}^{\prime}\right\}$ and $\left\{q_{n}\right\}$. The alignment parameters include: scale $\mathbf{s}$, rotation $\mathbf{R}$ and translation $\mathbf{t}$. These parameters can be obtained by the following optimization equation (Sanyal et al., 2019):
$Optim(\mathbf{s},\mathbf{R},\mathbf{t}=\min_{\mathbf{s},\mathbf{R},\mathbf{t}}\sum_{i}\left\|\mathbf{a}_{i}^{\prime}-\mathbf{s}\left(\mathbf{R}\cdot\mathbf{a}_{i}+\mathbf{t}\right)\right\|_{2}.$ (5)
@@ 4.4 Regularization Loss
To suppress the generation of distorted faces, we add the regularization loss which is commonly-used in face reconstruction task (Thies et al., 2016; Tewari et al., 2018; Deng et al., 2019; Shang et al., 2020):
@@
where $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, and $\boldsymbol{\gamma}$ are 3DMM parameter vectors that the network predicted. $\omega_{\alpha}$, $\omega_{\beta}$ and $\omega_{\gamma}$ are the weights for 3DMM parameter vectors. Following Deep3DFace (Deng et al., 2019), we set them to $1$, $0.8$ and $0.017$ with fine tuning.
@@ 4.5 Overall Loss
Note that 3D landmarks is not a must for our method. For general weakly supervised methods, we do not incorporate $L_{l\_3d}$ for comparison (Deng et al., 2019; Shang et al., 2020). Since semi-supervised methods HRN (Lei et al., 2023) adopts the entire 3D scans to attain training label, we add $L_{l\_3d}$ for comparative analysis.
@@ 5.1 Implementation Details
Dataset. Following BiseNet (Yu et al., 2018), we adopt CelebAMask-HQ (Lee et al., 2020) to pretrain Face Parse Net. Furthermore, Pixel-Face (Lyu et al., 2020) and Multi-PIE (Gross et al., 2008) are introduced to provide 3D landmarks label and multi-view faces. Pixel-Face contains $855$ subjects ranging in age from $18$ to $80$ years old. Each subject has $7$ or $23$ samples of different expressions. Pixel-Face has 3D mesh file of each sample as ground-truth but not 3DMM parameters or angle of multi-view images. In the experiment, the training/test split was set to $0.8$. For Multi-PIE, we leverage the version provided by TP-GAN (Huang et al., 2017), which will reduce a lot of data preprocessing work. This dataset contains about 140k multi-view images and does not provide 3D mesh and landmarks. We utilize the method of STNet (Chen et al., 2016) to detect and align faces. Then each input image is cropped and resized to 224$\times$224.
Network. Our network structure is shown in Fig. 2 and described in the methodology section. Based on the pre-trained BiseNet (Yu et al., 2018) with frozen weights, the face parse network is located in the beginning and end of the network. We implement DF-MVR by pytorch (Paszke et al., 2019). It is trained on 2 NVIDIA RTX 3090 GPUs card with a batch size of 16 for 512 epochs. We adopt Adam as the gradient descent method. The initial learning rate is set to 0.0001 which decays with the cosine anneal schedule. The average inference time of DF-MVR is 244ms.
Evaluation Metric. Following the previous works, RMSE (mm) (Wu et al., 2019; Deng et al., 2019; Shang et al., 2020) is used to compute point-to-plane L2 distance between predict 3D scans and ground-truth 3D scans. Concretely, the front face area is cropped for metrics calculation instead of using a complete BFM model (Sanyal et al., 2019; Deng et al., 2019; Shang et al., 2020). Before calculating point-to-plane L2 distance, the predicted 3D scans need to be registered with ground-truth 3D scans. Also, we take the ICP registration method (Li et al., 2017), the same as Deep3DFace (Deng et al., 2019).
@@ 5.2 Comparison to SOTAs
We compare our method with the existing weakly supervised MVRs on both Pixel-Face and Bosphorus datasets. More specifically, only the four methods with codes can be found in the literature related to multi-view weakly supervised 3D face reconstruction. MGCNet (Shang et al., 2020) takes multiple images for training, and then uses one single image for testing. We select the best results among the three images for display. Deep3DFace (Deng et al., 2019) does not release their source codes of its scoring network. Therefore, we use their codes to train/test on Pixel-Face to select the best results among the three images. Moreover, MVFR (Xiao et al., 2022) necessitates scale_RT data, rendering it currently effective only for the Facescape dataset (Yang et al., 2020; Zhu et al., 2023). HRN (Lei et al., 2023) does not release its training code, but it provides a pretrained model based on numerous data. To this end, we use its pretrained model and source code for multi-view face reconstruction. Given the scarcity of open sources for this task, we release our full codes (training & test) and pretrained model in GitHub [^4] https://github.com/weiguangzhao/DF_MVR.
@@
Deep3DFace (Deng et al., 2019) Pixel-Face 1.6641
MGCNet (Shang et al., 2020) Pixel-Face 1.8877
@@
Deep3DFace (Deng et al., 2019) Bosphorus 1.4777
MGCNet (Shang et al., 2020) Bosphorus 1.4418
Ours Bosphorus 1.3994
${}^{\dagger}$ (Lei et al., 2023) Pixel-Face 1.4061
@@
Comparsion on Pixel-Face. The quantitative results of the comparison are shown in Table 1. Since Deep3DFace (Deng et al., 2019) and MGCNet (Shang et al., 2020) did not use 3D landmarks, to be fair, we also provide the results of our model without using 3D landmarks for comparison. As observed, Our model (without 3D landmarks) shows $5.2\%$ improvement compared to the existing methods. While HRN (Lei et al., 2023) adopts the complete 3D ground-truth scan to attain deformation map and displacement map, we only take very light annotation, 101 3D landmarks to constraint the face reconstruction. ${}^{\dagger}$ shows very competitive results with fewer 3D annotation, compared to ${}^{\dagger}$.
@@
Deep3DFace (Deng et al., 2019) Pixel-Face 1.6641
MGCNet (Shang et al., 2020) Pixel-Face 1.8877
@@ 5.4 Ablation Study
In order to verify the effectiveness of MulEn-Unet and the mask mechanism we designed, we perform more ablation experiments as shown in Table 3. The mean and standard deviation of RMSE are again used as the evaluation metric. First, from v1, v2, v6, it can be found that the multi-view feature fusion network we designed is superior to traditional CNN and Unet in this task. Then, the results of v2 and v6 hint that the multi-layer feature interaction in the feature extraction stage is better than the direct concatenation of features at the end. To be fair, we set the number of layers of RedNet and ResNet to 50. Through the RMSE of v3 and v6, it is clear that RedNet performs better than ResNet in this task. For v4, we not only remove the mask loss but also the face mask $\mathbf{I}_{A}$, $\mathbf{I}_{B}$ and $\mathbf{I}_{C}$, which are concatenated to the original image. By comparing v5 and v6, we can see that the face mask mechanism promotes the network to generate a higher-precision model. Moreover, we verify the face mask and mask loss separately. Mask loss has a much smaller effect on the accuracy of the model than face mask. While the qualitative result of mask loss is positive for the adjustment of the model during training, we retain this part for reference. Finally, we remove $L_{lan\_3d}$, which means that our model can be trained with only three multi-view images without any 3D label (as denoted as v5). The result also shows that our model is accurate and stable.
@@
v3 MulEn-Unet ResNet Y Y 1.5206
v4 MulEn-Unet RedNet N Y 1.4656
v5 MulEn-Unet RedNet Y N 1.5770
v6 MulEn-Unet RedNet Y Y 1.4040
@@ 5.5 Robustness Analysis
Deep3DFace (Deng et al., 2019) 1.5030 1.4975 1.4924 1.4777 1.4905
MGCNet  (Shang et al., 2020) 1.4419 1.4412 1.4413 1.4418 1.4416
@@
We also validate the angle robustness of our model visually. These results are plotted in Fig. 8(a). Detailed error maps are introduced to display the error for each facial area directly. Take the second sample of Fig. 8(b) for a typical example where we fix $\mathbf{I}_{B}$ and $\mathbf{I}_{c}$ while changing the face angle of $\mathbf{I}_{A}$ from $10^{\circ}$ to $45^{\circ}$. Obviously, the error maps generated from these angles do not change much. Furthermore, we provide the error maps of Deep3DFace (Deng et al., 2019) for comparison. At different angles, DF-MVR reconstructs the face with less error, especially in critical areas (e.g. eye, brow, nose and mouth).